Amazing Manufacturing, Inc., has been considering the purchase of a new manufacturing facility for $580,000. The facility is to be fully depreciated on a straight-line basis over seven years. It is expected to have no resale value at that time. Operating revenues from the facility are expected to be $430,000, in nominal terms, at the end of the first year. The revenues are expected to increase at the inflation rate of 4 percent. Production costs at the end of the first year will be $275,000, in nominal terms, and they are expected to increase at 5 percent per year. The real discount rate is 7 percent. The corporate tax rate is 24 percent.

   

Calculate the NPV of the project. (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
Answer #1

Given

Purchase cost of new manufacturing facility = $580,000

Operation revenues = $430,000 in year1 with 4% increase every year

Production cost = $275,000 in year1 with 5% increase every year

The above given cashflows are nominal cashflows

But the discount rate of 7% is real discount rate. So, to calculate NPV we convert the real discount rate into nominal discount rate by using the formula

(1+real discount rate)\times(1+inflation rate) = (1+nominal discount rate)

(1+0.07)\times(1+0.04) = (1.1128) = (1+nominal discount rate)

Nominal discount rate = 1.1128-1 = 0.1128

\therefore Nominal discount rate = 11.28%

Depreciation per annum = $580,000 \div 7years =$82857.14

Computation of NPV

Particulars Year 1 Year 2 Year3 Year 4 Year 5 Year 6 Year 7
Revenues $430,000 $447,200 $465,088 $483,691.52 $503,039.1808 $523,160.748 $544,087.1779
Production costs $275,000 $288,750 $303,187.5 $318,346.875 $334,264.2187 $350,977.4296 $368,526.3011
Depreciation $82,857.14 $82,857.14 $82,857.14 $82,857.14 $82,857.14 $82,857.14 $82,857.16
Profit before Tax $72,142.86 $75,592.86 $79,043.36 $82,487.505 $85,917.8221 $89,326.1784 $92,703.7168
Tax @ 24% $17,314.2864 $18,142.2864 $18,970.4064 $19,797.0012 $20,620.2773 $21,438.2828 $22,248.8920
Profit after Tax $54,828.5736 $57,450.5736 $60,072.9536 $62,690.5038 $65,297.5448 $67,887.8956 $70,454.8248
Add: Depreciation $82,857.14 $82,857.14 $82,857.14 $82,857.14 $82,857.14 $82,857.14 $82,857.16
Cashflows after tax $137,685.7136 $140,307.7136 $142,930.0936 $145,547.6438 $148,154.6848 $150,745.0356 $153,311.9648
PvF(11.28%,n) 0.899 0.807 0.726 0.652 0.586 0.527 0.473
Present value of cash inflows $123,779.456 $113,228.3248 $103,767.2479 $94,897.0638 $86,818.6452 $79,442.6338 $72,516.5593

NPV = Present value of cash inflows - Present value of cash outflows

= $674,449.9308 - $580,000 = $94,449.93

Since NPV is positive, company can take up the new manufacturing facility